OpenStudy (msmr):

A sprinter can burst out of the blocks to her top speed of 11.5 m/s in the first 5.0 m of the race. How long does it take her to reach her top speed?

OpenStudy (theeric):

Hello! The sprinter is gaining velocity, and so the sprinter is changing velocity. That is acceleration! Do you know the definition of acceleration?

OpenStudy (theeric):

I'd use \(d=d_0+v_0\ t+\frac{1}{2}a\ t^2\). Here, \(d_0\) and \(v_0\) are \(0\), so \(d=\frac{1}{2}a\ t^2\). And I asked you about the definition of acceleration: change in velocity per change in time. \(a=\dfrac{v-v_0}{t-t_0}=\dfrac{v}{t}\) Putting that into the equation before it (substituting \(a\)), what do you get? And can you solve it? Good luck! :)

OpenStudy (msmr):

I got 2.6 seconds; is that correct?

OpenStudy (theeric):

It might be, but I got a different answer..

OpenStudy (msmr):

what did you get?

OpenStudy (theeric):

What equation did you get?

OpenStudy (theeric):

I'm still pretty sure that it is not 2.6 s. What work did you do? I can help get you on track, if you're off.

OpenStudy (amistre64):

since distance (displacement) is equal to the area under the velocity curve .... determine the value for t such that the area of the triangle is 5